MICHIGAN STATE UNIVERSITY Department of Mathematics HW08: Problem 6 Problem 2 40 45 50 Mileage (in MPG 60 65 7becane
media%2F3f2%2F3f2b4fee-72ea-4a08-bd8a-74
Answer #1

NULL HYPOTHESIS H0: μ = 50 MPG

ALTERNATIVE HYPOTHESIS Ha: \mu\ne50 MPG (TWO TAILED TEST)

t= (\overline{x}-\mu)/s/\sqrt(n)

t= (53.3-50)/5.2/\sqrt(14)

t= 3.3/5.2/3.74

t= 3.3/1.39

t= 2.374

Degrees of freedom= 14-1=13

The P-Value is .034. The result is significant because p < .05.

Decison: REJECT NULL HYPOTHESIS H0

95% CONFIDENCE INTERVAL

t CRITICAL = 2.16
sM = √(5.2^2/14) = 1.39

μ = M ± t(sM)
μ = 53.3 ± 2.16*1.39
μ = 53.3 ± 3.002 95% CI [50.298, 56.302].

You can be 95% confident that the population mean (μ) falls between 50.298 and 56.302.